OpenStudy (anonymous):

Can you factor 2t^2+t-6

OpenStudy (anonymous):

(2t-3)(t+2)

OpenStudy (dumbcow):

multiply 2*-6 = -12 4* -3 = -12 4+ -3 = 1 (2t-3)(t+2)

OpenStudy (anonymous):

wow I'm a dunce.

OpenStudy (anonymous):

(2t - 3)(t + 2)

OpenStudy (anonymous):

\[2t^{2}+t-6 = 2t^{2}+4t-3t-6 = 2t(t+2)-3(t+2) = (2t-3)(t+2) \]